### v4

@@ 3.2 Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain (Sec. 3.2.1), Time Domain (Sec. 3.2.2), Frequency Domain (Sec. 3.2.3), and Data-Driven (Sec. 3.2.4). We also summarize the detailed information about popular methods in Table 6.
3.2.1 Space Domain
@@
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color he2019detection , saturation mccloskey2019detecting , artifacts zhao2021multi ; shiohara2022detecting ; cao2022end , gradient variations tan2023learning , etc. Specifically, RECCE cao2022end considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad tan2023learning utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions nirkin2021deepfake , as well as the fine-grained details of image textures chai2020makes ; liu2020global . Recently, Ba et alba2024exposing focuse not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
$\bullet$ Local Noise Inconsistency. Image forgery may involve adding, modifying, or removing content in the image, potentially altering the noise distribution in the image. Detection methods based on noise aim to identify such local or even global differences in the image. Zhou et alzhou2017two propose a dual-stream structure, combining GoogleNet with a triplet network to focus on tampering artifacts and local noise in images. Nguyen et alnguyen2019capsule utilize capsule networks to detect forged pictures and videos in various forgery scenarios. NoiseDF wang2023noise specializes in identifying underlying noise traces left behind in Deepfake videos, introducing an efficient and novel Multi-Head Relative Interaction with depth-wise separable convolutions to enhance detection performance.
3.2.2 Time Domain
@@
$\bullet$ Abnormal Physiological Information. Forgery videos often overlook the authentic physiological features of humans, failing to achieve overall consistency with authentic individuals. Therefore, some methods focus on assessing the plausibility of the physiological features of the generated faces in videos. Li et alli2018ictu detect blinking and blink frequency in videos as criteria for determining the video’s authenticity. Yang et alyang2019exposing focuses on the inconsistency of head poses in videos, comparing the differences between head poses estimated using all facial landmarks and those estimated using only the landmarks in the central region. Peng et alpeng2024deepfakes focuse on inter-frame gaze angles, obtaining gaze characteristics of each video frame and using a spatio-temporal feature aggregator to combine temporal gaze features, spatial attribute features, and spatial texture features as the basis for detection and classification.
@@
$\bullet$ Inter-Frame Inconsistency. Methods zheng2021exploring ; gu2022delving ; yin2023dynamic ; yang2023masked ; choi2024exploiting ; xu2024towards based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et algu2022delving focuse on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et alyin2023dynamic design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et alyang2023masked approach forgery detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et alchoi2024exploiting discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
@@
$\bullet$ Multimodal Inconsistency. The core idea behind multimodal detection algorithms is to make judgments based on the flow of prior information from multiple attributes rather than solely considering the image or audio differences of individual characteristics in each frame. The consideration of audio-visual modal inconsistency has received extensive research in various methods haliassos2022leveraging ; cozzolino2023audio ; feng2023self ; yang2023avoid . POI-Forensics cozzolino2023audio proposes a deep forgery detection method based on audio-visual authentication, utilizing contrastive learning to learn the most distinctive embeddings for each identity in moving facial and audio segments. AVoiD-DF yang2023avoid embeds spatiotemporal information in a spatiotemporal encoder and employs a multimodal joint decoder to fuse multimodal features and learn their inherent relationships. Subsequently, a cross-modal classifier is applied to detect disharmonious operations within and between modalities. Agarwal et alagarwal2021detecting describe a forensic technique for detecting fake faces using static and dynamic auditory ear characteristics. Indeed, multimodal detection methods are currently a hotspot in forgery detection research.
3.2.3 Frequency Domain
@@
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works qian2020thinking ; frank2020leveraging ; li2021frequency ; miao2022hierarchical ; tan2024frequency utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, F3-Net qian2020thinking proposes a dual-branch framework. One frequency-aware branch utilizes Frequency-aware Image Decomposition (FAD) to learn subtle forgery patterns in suspicious images. In contrast, the other branch aims to extract high-level semantics from Local Frequency Statistics (LFS) to describe the frequency-aware statistical differences between real and forged faces. HFI-Net miao2022hierarchical consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et altan2024frequency introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration masi2020two ; guo2023constructing , and Guo et alguo2023constructing design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
3.2.4 Data Driven
@@
Data-driven forgery detection focuses on learning specific patterns and features from extensive image or video datasets to distinguish between genuine and potentially manipulated images. Some methods hsu2018learning ; yu2019attributing believe that images generated by specific models possess unique model fingerprints. Based on this belief, forgery detection can be achieved by focusing on the model’s training. In addition, FakeSpotter wang2021fakespotter introduces the Neuron Coverage Criterion to capture layer-wise neuron activation behavior. It monitors the neural behavior of a deep face recognition system through a binary classifier to detect fake faces. There are also methods zhao2021learning ; huang2023implicit that attempt to classify the sources of different components in an image. For instance, Huang et alhuang2023implicit think that the difference between explicit and implicit identity helps detect face swapping. There are numerous data-driven methods dang2020detection ; hu2022finfer ; guo2023hierarchical ; guo2023controllable , and it is not feasible to discuss each one in detail here.
3.3 Specific Related Domains
In this section, we briefly review related popular tasks beyond the deepfake generation, such as Face Super-resolution (Sec. 3.3.1), Portrait Style Transfer (Sec. 3.3.2), Body Animation (Sec. 3.3.3), and Makeup Transfer (Sec. 3.3.4).
3.3.1 Face Super-resolution
@@
$\bullet$ Convolutional Neural Networks. Early works chen2019sequential ; hu2019face ; lu2020global on facial super-resolution based on CNNs aims to leverage the powerful representational capabilities of CNNs to learn the mapping relationship between low-resolution and high-resolution images from training samples. Depending on whether they focus on local details of the image, they can be divided into global methods chen2019sequential ; huang2017wavelet , local methods hu2019face , and mixed methods lu2020global .
@@
$\bullet$ Generative Adversarial Network. GAN aims to achieve the optimal output result through an adversarial process between the generator and the discriminator. This type of method zhang2019ranksrgan ; pan2024lpsrgan currently dominates the field for flexible and efficient architecture.
3.3.2 Portrait Style Transfer
@@
$\bullet$ Generative Adversarial Network. The most mature style transfer algorithm is the GAN-based approach abdal20233davatargan ; xu2022your3demoji ; jiang2023scenimefy . However, due to the relatively poor stability of GANs, it is common for the generated images to contain artifacts and unreasonable components. 3DAvatarGAN abdal20233davatargan bridges the pre-trained 3D-GAN in the source domain with the 2D-GAN trained on an artistic dataset to achieve cross-domain generation. Scenimefy jiang2023scenimefy utilizes semantic constraints provided by text models like CLIP to guide StyleGAN generation and applies patch-based contrastive style loss to enhance stylization and fine details further.
@@
$\bullet$ Diffusion-based methods liu2023portrait ; hur2024expanding ; lei2023diffusiongan3d represent the generative process of cross-domain image transfer using diffusion processes. DiffusionGAN3D lei2023diffusiongan3d combines 3DGAN chan2022efficient with a diffusion model from text to graphics, introducing relative distance loss and learnable tri-planes for specific scenarios to further enhance cross-domain transformation accuracy.
3.3.3 Body Animation
@@
$\bullet$ Generative Adversarial Network. GAN-based approaches zhao2022thin ; jain2023vgflow ; zhou2022cross aim to train a model to generate images whose conditional distribution resembles the target domain, thus transferring information from reference images to target images. CASD zhou2022cross is based on a style distribution module using a cross-attention mechanism, facilitating pose transfer between source semantic styles and target poses. VGFlow jain2023vgflow introduces a visibility-guided flow module to preserve texture and perform style manipulation concurrently. However, existing methods still rely considerably on training samples, and exhibit decreased performance when dealing with actions in rare poses.
@@
$\bullet$ Diffusion-based. The task of body animation using diffusion models aims to utilize diffusion processes to generate the propagation and interaction of movements between body parts based on a reference source. This approach wang2023leo ; ma2023follow ; hu2023animate represents a current hot topic in research and implementation. LEO wang2023leo focuses on the spatiotemporal continuity between generated actions, employing the Latent Motion Diffusion Model to represent motion as a series of flow graphs during the generation process. Animate Anyone hu2023animate harnesses the powerful generation capabilities of stable diffusion models combined with attention mechanisms to generate high-quality character animation video.
3.3.4 Makeup Transfer
@@
$\bullet$ Graphics-based Approaches. Before the introduction of neural networks, traditional computer graphics methods guo2009digital ; li2015simulating use image gradient editing and physics-based operations to understand the semantics of makeup. By decomposing the input image into multiple layers, each representing different facial information, traditional methods would distort the reference facial image onto the non-makeup target using facial landmarks for each layer. However, due to the limitations of manually designed operators, the output images from traditional methods often appear unnatural, with noticeable artifacts. Additionally, there is a tendency for background information to be modified to some extent.
@@
$\bullet$ Generative Adversarial Network. Early deep learning-based methods gu2019ladn aim at fully automatic makeup transfer. However, these methods chang2018pairedcyclegan ; chen2019beautyglow exhibit poor performance when faced with significant differences in pose and expression between the source and target faces and are unable to handle extreme makeup scenarios well. Some methods jiang2020psgan ; liu2021psgan++ ; zhong2023sara proposes their solutions, PSGAN++ liu2021psgan++ comprises the Makeup Distillation Network, Attentive Makeup Morphing module, Style Transfer Network, and Identity Extraction Network, further enhancing the ability of PSGAN jiang2020psgan to perform targeted makeup transfer with detail preservation. ELeGANt yang2022elegant , CUMTGAN hao2022cumtgan , and HT-ASE li2023hybrid explore the preservation of detailed information. ELeGANt yang2022elegant encodes facial attributes into pyramid feature maps to retain high-frequency information. Matching facial semantic information, which involves rendering makeup styles onto semantically corresponding positions of a target image, is often overlooked. SSAT sun2022ssat introduces the SSCFT module and weakly supervised semantic loss for accurate semantic correspondence. SSAT++ sun2023ssat further improves color fidelity matching, but both models are complex with high training complexity. In addition, BeautyREC yan2023beautyrec , based on the Transformer with long-range visual dependencies, achieves efficient global makeup transfer with significantly reduced overall model parameters compared to previous works.
4 Benchmark Results
We firstly introduce the evaluation metrics commonly used for each deepfake tasks, and then we evaluate the performance of representative methods for each reviewed field on the most widely used datasets with the data sourced from the original respective papers. Considering the differences in training datasets, testing datasets, and metrics used by different approaches, we strive to compare them fairly in each table to the greatest extent.
4.1 Metrics
@@
$\bullet$ Face Swapping. The most commonly used objective evaluation metrics for face swapping include ID Ret, Expression Error, Pose Error, and FID. ID Ret is calculated by a pre-trained face recognition model wang2018cosface , measuring the Euclidean distance between the generated face and the source face. A higher ID Ret indicates better preservation of identity information. Expression and pose errors quantify the differences in expression and pose between the generated face and the source face. These metrics are evaluated using a pose estimator ruiz2018fine and a 3D facial model deng2019accurate , extracting expression and pose vectors for the generated and source faces. Lower values for expression error and pose error indicate higher facial expression and pose similarity between the swapped face and the source face. FID  heusel2017gans is used to assess image quality, with lower FID values indicating that the generated images closely resemble authentic facial images in appearance.
@@
$\bullet$ Face Reenactment. Face reenactment commonly uses consistent evaluation metrics, including CSIM, SSIM wang2004image , PSNR, LPIPS zhang2018unreasonable , LMD chen2018lip , and FID. CSIM describes the cosine similarity between the generated and source faces, calculated by ArcFace deng2019arcface , with higher values indicating better performance. SSIM, PSNR, LPIPS, and FID are used to measure the quality of synthesized images. SSIM measures the structural similarity between two images, with higher values indicating a closer resemblance to natural images. PSNR quantifies the ratio between a signal’s maximum possible power and noise’s power, indicating higher quality for higher values. LPIPS assesses reconstruction fidelity using a pre-trained AlexNet krizhevsky2012imagenet to extract feature maps for similarity score computation. As mentioned earlier, FID is used to evaluate image quality. LMD assesses the accuracy of lip shape in generated images or videos, with lower values indicating better model performance.
@@
$\bullet$ Talking Face Generation. Expanding upon face reenactment metrics, talking face generation incorporates additional metrics, including M/F-LMD, Sync, LSE-C, and LSE-D. The LSE-C and LSE-D are usually used to measure lip synchronization effectiveness prajwal2020lip . The landmark distances on the mouth (M-LMD) chen2019hierarchical and the confidence score of SyncNet (Sync) measure synchronization between the generated lip motion and the input audio. F-LMD computes the difference in the average distance of all landmarks between predictions and ground truth (GT) as a measure to assess the generated expression. In addition, there are some meaningful metrics, such as LSE-C and LSE-D for measuring lip synchronization effectiveness prajwal2020lip , AVD li2023one for evaluating identity preservation performance, AUCON ha2020marionette for assessing facial pose and expression jointly, and AGD doukas2023free for evaluating eye gaze changes. These newly proposed evaluation metrics enrich the performance assessment system by targeting various aspects of the model’s performance.
@@
$\bullet$ Facial Attribute Editing. The standard evaluation metrics used in face attribute manipulation are FID, LPIPS zhang2018unreasonable , KID binkowski2018demystifying , PSNR and SSIM. KID is one of the image quality assessment metrics commonly used in face editing work and other generative modeling tasks to quantify the difference in distribution between the generated image and the actual image, with lower KID values indicating better model performance. Some text-guide work will also use the CLIP Score to measure the consistency between the output image and the text, calculated as the cosine similarity between the normalized image and the text embedding. Higher values of CLIP Score indicate better consistency of the generated image with the corresponding text sentence.
4.2 Main Results on Deepfake Generation
We select widely used metrics for each task and tabulate the performance for representative methods.
$\bullet$ Results on Face Swapping. Table 7 displays the performance evaluation results of some representative models on the Face Swapping task using the FF++ rossler2019faceforensics++ dataset. WSC-Swapren2023reinforced captures facial region external attribute information and internal identity information through two independent encoders, which achieves good identity preservation and facial pose retention. However, it exhibits sub-optimal performance in facial expression error metrics.
$\bullet$ Results on Face Reenactment. Table 8 and Table 9 show the performance evaluation results on the VoxCeleb nagrani2017voxceleb dataset for self-reenactment and cross-subject reenactment, respectively. AVFR-GAN agarwal2023audio achieves better performance by using the multimodal modeling. 10 presents quality assessment results on the VoxCeleb2 dataset. HiDe-NeRF li2023one represents 3D scenes using canonical appearance fields and implicit deformation fields. It achieves accurate facial attribute modeling by explicitly decoupling facial pose and expression attributes using the deformation module.

### v5

@@ 3.2. Forgery Detection
In this section, we review current forgery detection techniques based on the type of detection cues, categorizing them into three: Space Domain, Time Domain, Frequency Domain, and Data-Driven. We also summarize the detailed information about popular methods in Tab. 6.
3.2.1. Space Domain
@@
$\bullet$ Image-level Inconsistency. The generation process of forged images often involves partial alterations rather than global generation, leading to common local differences in non-globally generated forgery methods. Therefore, some methods focus on differences in image spatial details as criteria for determining whether an image is forged, such as color (he2019detection, ), saturation (mccloskey2019detecting, ), artifacts (zhao2021multi, ; shiohara2022detecting, ; cao2022end, ), gradient variations (tan2023learning, ), etc. Specifically, RECCE (cao2022end, ) considers shadow generation from a training perspective, utilizing the learned representations on actual samples to identify image reconstruction differences. LGrad (tan2023learning, ) utilizes a pre-trained transformation model, converting images to gradients to visualize general artifacts and subsequently classifying based on these representations. In addition, some works focus on detection based on differences in facial and non-facial regions (nirkin2021deepfake, ), as well as the fine-grained details of image textures (chai2020makes, ; liu2020global, ). Recently, Ba et al(ba2024exposing, ) focuse not only on the discordance in a single image region but also on the detection of fused local representation information from multiple non-overlapping areas.
[…]
3.2.2. Time Domain
@@
$\bullet$ Abnormal Physiological Information. Forgery videos often overlook the authentic physiological features of humans, failing to achieve overall consistency with authentic individuals. Therefore, some methods focus on assessing the plausibility of the physiological features of the generated faces in videos. Li et al(li2018ictu, ) detect blinking and blink frequency in videos as criteria for determining the video’s authenticity. Yang et al(yang2019exposing, ) focuses on the inconsistency of head poses in videos, comparing the differences between head poses estimated using all facial landmarks and those estimated using only the landmarks in the central region. Peng et al(peng2024deepfakes, ) focuse on inter-frame gaze angles, obtaining gaze characteristics of each video frame and using a spatio-temporal feature aggregator to combine temporal gaze features, spatial attribute features, and spatial texture features as the basis for detection and classification.
@@
$\bullet$ Inter-Frame Inconsistency. Methods (zheng2021exploring, ; gu2022delving, ; yin2023dynamic, ; yang2023masked, ; choi2024exploiting, ; xu2024towards, ) based on inter-frame inconsistency for forgery detection aim to uncover differences in images between adjacent frames or frames with specific temporal spans. Gu et al(gu2022delving, ) focuse on inter-frame image inconsistency by densely sampling adjacent frames, while Yin et al(yin2023dynamic, ) design a Dynamic Fine-grained Difference Capturing module and a Multi-Scale Spatio-Temporal Aggregation module to cooperatively model spatio-temporal inconsistencies. Yang et al(yang2023masked, ) approach forgery detection as a graph classification problem, emphasizing the relationship information between facial regions to capture the relationships among local features across different frames. Choi et al(choi2024exploiting, ) discover that the style variables in each frame of Deepfake work change. Based on this, they developed a style attention module to focus on the inconsistency of the style latent variables between frames.
@@
$\bullet$ Multimodal Inconsistency. The core idea behind multimodal detection algorithms is to make judgments based on the flow of prior information from multiple attributes rather than solely considering the image or audio differences of individual characteristics in each frame. The consideration of audio-visual modal inconsistency has received extensive research in various methods (haliassos2022leveraging, ; cozzolino2023audio, ; yang2023avoid, ). POI-Forensics (cozzolino2023audio, ) proposes a deep forgery detection method based on audio-visual authentication, utilizing contrastive learning to learn the most distinctive embeddings for each identity in moving facial and audio segments. AVoiD-DF (yang2023avoid, ) embeds spatiotemporal information in a spatiotemporal encoder and employs a multimodal joint decoder to fuse multimodal features and learn their inherent relationships. Subsequently, a cross-modal classifier is applied to detect disharmonious operations within and between modalities. Agarwal et al(agarwal2021detecting, ) describe a forensic technique for detecting fake faces using static and dynamic auditory ear characteristics. Indeed, multimodal detection methods are currently a hotspot in forgery detection research.
3.2.3. Frequency Domain
@@
Frequency domain-based forgery detection methods transform image time-domain information into the frequency domain. Works (qian2020thinking, ; frank2020leveraging, ; li2021frequency, ; miao2022hierarchical, ; tan2024frequency, ) utilize statistical measures of periodic features, frequency components, and frequency characteristic distributions, either globally or in local regions, as evaluation metrics for forgery detection. Specifically, F3-Net (qian2020thinking, ) proposes a dual-branch framework. One frequency-aware branch utilizes Frequency-aware Image Decomposition (FAD) to learn subtle forgery patterns in suspicious images. In contrast, the other branch aims to extract high-level semantics from Local Frequency Statistics (LFS) to describe the frequency-aware statistical differences between real and forged faces. HFI-Net (miao2022hierarchical, ) consists of a dual-branch network and four Global-Local Interaction (GLI) modules. It effectively explores multi-level frequency artifacts, obtaining frequency-related forgery clues for face detection. Tan et al(tan2024frequency, ) introduce a novel frequency-aware approach called FreqNet, which focuses on the high-frequency information of images and combines it with a frequency-domain learning module to learn source-independent features. Furthermore, some approaches combine spatial, temporal, and frequency domains for joint consideration (masi2020two, ; guo2023constructing, ), and Guo et al(guo2023constructing, ) design a spatial-frequency interaction convolution to construct a novel backbone network for Deepfake detection.
3.2.4. Data Driven
@@
Data-driven forgery detection (dang2020detection, ; hu2022finfer, ; guo2023hierarchical, ; guo2023controllable, ) focuses on learning specific patterns and features from extensive image or video datasets to distinguish between genuine and potentially manipulated images. Some methods (hsu2018learning, ; wang2021fakespotter, ) believe that images generated by specific models possess unique model fingerprints. Based on this belief, forgery detection can be achieved by focusing on the model’s training. In addition, FakeSpotter (wang2021fakespotter, ) introduces the Neuron Coverage Criterion to capture layer-wise neuron activation behavior. It monitors the neural behavior of a deep face recognition system through a binary classifier to detect fake faces. There are also methods (zhao2021learning, ; huang2023implicit, ) that attempt to classify the sources of different components in an image. For instance, Huang et al(huang2023implicit, ) think that the difference between explicit and implicit identity helps detect face swapping.
3.3. Specific Related Domains
3.3.1. Face Super-resolution
@@
$\bullet$ Convolutional Neural Networks. Early works (chen2019sequential, ; hu2019face, ; lu2020global, ) on facial super-resolution based on CNNs aims to leverage the powerful representational capabilities of CNNs to learn the mapping relationship between low-resolution and high-resolution images from training samples. Depending on whether they focus on local details of the image, they can be divided into global methods (chen2019sequential, ; huang2017wavelet, ), local methods (hu2019face, ), and mixed methods (lu2020global, ).
@@
$\bullet$ Generative Adversarial Network. GAN aims to achieve the optimal output result through an adversarial process between the generator and the discriminator. This type of method (zhang2019ranksrgan, ; pan2024lpsrgan, ) currently dominates the field for flexible and efficient architecture.
3.3.2. Portrait Style Transfer
@@
$\bullet$ Generative Adversarial Network. The most mature style transfer algorithm is the GAN-based approach (abdal20233davatargan, ; xu2022your3demoji, ; jiang2023scenimefy, ). However, due to the relatively poor stability of GANs, it is common for the generated images to contain artifacts and unreasonable components. 3DAvatarGAN (abdal20233davatargan, ) bridges the pre-trained 3D-GAN in the source domain with the 2D-GAN trained on an artistic dataset to achieve cross-domain generation. Scenimefy (jiang2023scenimefy, ) utilizes semantic constraints provided by text models like CLIP to guide StyleGAN generation and applies patch-based contrastive style loss to enhance stylization and fine details further.
@@
$\bullet$ Diffusion-based. Diffusion-based methods (liu2023portrait, ; hur2024expanding, ; lei2023diffusiongan3d, ) represent the generative process of cross-domain image transfer using diffusion processes. DiffusionGAN3D (lei2023diffusiongan3d, ) combines 3DGAN with a diffusion model from text to graphics, introducing relative distance loss and learnable tri-planes for specific scenarios to further enhance cross-domain transformation accuracy.
3.3.3. Body Animation
@@
$\bullet$ Generative Adversarial Network. GAN-based approaches (zhao2022thin, ; zhou2022cross, ) aim to train a model to generate images whose conditional distribution resembles the target domain, thus transferring information from reference images to target images. CASD (zhou2022cross, ) is based on a style distribution module using a cross-attention mechanism, facilitating pose transfer between source semantic styles and target poses. However, existing methods still rely considerably on training samples, and exhibit decreased performance when dealing with actions in rare poses.
@@
$\bullet$ Diffusion-based. The task of body animation using diffusion models aims to utilize diffusion processes to generate the propagation and interaction of movements between body parts based on a reference source. This approach (wang2023leo, ; ma2023follow, ) represents a current hot topic in research and implementation. LEO (wang2023leo, ) focuses on the spatiotemporal continuity between generated actions, employing the Latent Motion Diffusion Model to represent motion as a series of flow graphs during the generation process.
3.3.4. Makeup Transfer
@@
$\bullet$ Graphics-based Approaches. Before the advent of neural networks, traditional computer graphics methods (guo2009digital, ; li2015simulating, ) relied on image‐gradient editing and physics-based operations to interpret makeup semantics. These approaches decomposed an input image into multiple layers and warped the reference makeup image onto the target using facial landmarks. However, the reliance on manually designed operators often led to unnatural results, including visible artifacts and unintended alterations to background regions.
@@
$\bullet$ Generative Adversarial Network. Early deep learning-based methods (gu2019ladn, ) aim at fully automatic makeup transfer. However, these methods (chang2018pairedcyclegan, ; chen2019beautyglow, ) exhibit poor performance when faced with significant differences in pose and expression between the source and target faces and are unable to handle extreme makeup scenarios well. Some methods (jiang2020psgan, ; liu2021psgan++, ; zhong2023sara, ) proposes their solutions, PSGAN++ (liu2021psgan++, ) comprises the Makeup Distillation Network, Attentive Makeup Morphing module, Style Transfer Network, and Identity Extraction Network, further enhancing the ability of PSGAN (jiang2020psgan, ) to perform targeted makeup transfer with detail preservation. ELeGANt (yang2022elegant, ), CUMTGAN (hao2022cumtgan, ), and HT-ASE (li2023hybrid, ) explore the preservation of detailed information.
4. Benchmark
We introduce the evaluation metrics commonly used for each deepfake task, followed by a summary of the performance of representative methods on widely adopted datasets based on the results reported in their original papers. Given the variations in testing setups, and evaluation criteria across different approaches, we strive to present fair and consistent comparisons in each table.
4.1. Metrics
@@
$\bullet$ Face Swapping. The most commonly used objective evaluation metrics for face swapping include ID Ret, Expression Error, Pose Error, and FID. ID Ret is calculated by a pre-trained face recognition model (wang2018cosface, ), measuring the Euclidean distance between the generated face and the source face. A higher ID Ret indicates better preservation of identity information. Expression and pose errors quantify the differences in expression and pose between the generated face and the source face. These metrics are evaluated using a pose estimator (ruiz2018fine, ) and a 3D facial model (deng2019accurate, ), extracting expression and pose vectors for the generated and source faces. Lower values for expression error and pose error indicate higher facial expression and pose similarity between the swapped face and the source face. FID  (heusel2017gans, ) is used to assess image quality, with lower FID values indicating that the generated images closely resemble authentic facial images in appearance.
@@
$\bullet$ Face Reenactment. Face reenactment commonly uses consistent evaluation metrics, including CSIM, SSIM (wang2004image, ), PSNR, LPIPS (zhang2018unreasonable, ), LMD (chen2018lip, ), and FID. CSIM describes the cosine similarity between the generated and source faces, calculated by ArcFace (deng2019arcface, ), with higher values indicating better performance. SSIM, PSNR, LPIPS, and FID are used to measure the quality of synthesized images. SSIM measures the structural similarity between two images, with higher values indicating a closer resemblance to natural images. PSNR quantifies the ratio between a signal’s maximum possible power and noise’s power, indicating higher quality for higher values. LPIPS assesses reconstruction fidelity using a pre-trained AlexNet (krizhevsky2012imagenet, ) to extract feature maps for similarity score computation. As mentioned earlier, FID is used to evaluate image quality. LMD assesses the accuracy of lip shape in generated images or videos, with lower values indicating better model performance.
@@
$\bullet$ Talking Face Generation. Expanding upon face reenactment metrics, talking face generation incorporates additional metrics, including M/F-LMD, Sync, LSE-C, and LSE-D. The LSE-C and LSE-D are usually used to measure lip synchronization effectiveness (prajwal2020lip, ). The landmark distances on the mouth (M-LMD) (chen2019hierarchical, ) and the confidence score of SyncNet (Sync) measure synchronization between the generated lip motion and the input audio. F-LMD computes the difference in the average distance of all landmarks between predictions and ground truth (GT) as a measure to assess the generated expression. In addition, there are some meaningful metrics, such as LSE-C and LSE-D for measuring lip synchronization effectiveness (prajwal2020lip, ), AVD (li2023one, ) for evaluating identity preservation performance, AUCON (ha2020marionette, ) for assessing facial pose and expression jointly, and AGD (doukas2023free, ) for evaluating eye gaze changes. These newly proposed evaluation metrics enrich the performance assessment system by targeting various aspects of the model’s performance.
@@
$\bullet$ Facial Attribute Editing. The standard evaluation metrics used in face attribute manipulation are FID, LPIPS (zhang2018unreasonable, ), KID (binkowski2018demystifying, ), PSNR and SSIM. KID is one of the image quality assessment metrics commonly used in face editing work and other generative modeling tasks to quantify the difference in distribution between the generated image and the actual image, with lower KID values indicating better model performance. Some text-guide work will also use the CLIP Score to measure the consistency between the output image and the text, calculated as the cosine similarity between the normalized image and the text embedding. Higher values of CLIP Score indicate better consistency of the generated image with the corresponding text sentence.
4.2. Benchmark Protocol
In the absence of a unified benchmark, this survey adopts commonly used datasets and evaluation metrics for each task to establish a reference benchmarking protocol. All reported results are drawn directly from the original publications; therefore, the benchmark is intended for performance presentation rather than strict comparative evaluation.
$\bullet$ Face Swapping. This task is primarily evaluated on the FF++ (rossler2019faceforensics++, ) dataset. A test set is constructed by uniformly sampling 10 frames from each of 1,000 videos (10,000 images in total), and performance is measured using ID Retention, Expression Error, Pose Error, and FID. However, differences in training datasets may affect result comparability. In this survey, we follow this evaluation protocol and explicitly report experimental settings that may impact fairness.  Tab. 8 summarizes the results reported in the original publications and documents training data differences. Notably, RAFSwap (xu2022region, ) and Xu et al(xu2022high, ) adopted the MegaFS (zhu2021one, ) preprocessing protocol for FF++, while SimSwap (chen2020simswap, ) and CanonSwap (luo2025canonswap, ) applied pixel-threshold filtering to improve data quality.
$\bullet$ Face Reenactment. This task is evaluated under three settings: self-reenactment, cross-identity reenactment, and quality assessment. VoxCeleb (nagrani2017voxceleb, ) is used for self- and cross-identity reenactment, while VoxCeleb2 (chung2018voxceleb2, ) is adopted for quality assessment. For self-reenactment, the first frame of each VoxCeleb test video serves as the source image and the remaining frames as driving images. For cross-identity reenactment, 35 video pairs with different identities are randomly selected from the VoxCeleb test set. Self-reenactment is evaluated using CSIM, PSNR, LPIPS, FID, and SSIM; cross-identity reenactment uses CSIM, AVD, AUCON, FID, and AGD; and quality assessment adopts CSIM, PSNR, LMD, FID, and SSIM. Tab. 810 report the original results and document differences in training datasets.
$\bullet$ Talking Face Generation. Since the MEAD (wang2020mead, ) dataset provides explicit emotion annotations, it is commonly adopted as the benchmark for this task. As shown in  Tab. 12, this survey follows this protocol and uses MEAD as the benchmark dataset. The evaluation metrics include CSIM, LMD, M/F-LMD, Sync, FID, PSNR, and SSIM. We report the results from the original publications of the selected methods and document their corresponding training dataset configurations. It is worth noting that the selected methods differ in their use of the MEAD dataset, and the validation set splits are not standardized across studies.
$\bullet$ Facial Attribute Editing. This task is typically evaluated in terms of facial reconstruction capability and FID, yet lacks a unified evaluation protocol. Therefore, as shown in  Tab. 12 and  Tab. 14, this survey does not impose fixed training or validation splits; instead, it reports representative methods’ PSNR, LPIPS, SSIM, and FID results as presented in the original publications.